# Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks

### v1

@@ 5 A stable equivalence of Morita type
Denote $\hat{E}_{\mathfrak{h}}/(j+J(\mathbb{A}_{\delta})^{D})$ by $\bar{\hat{E}}_{\mathfrak{h}}$. Then for any $\mathfrak{a}\in\hat{E}_{\mathfrak{h}}$, denoting by $\bar{\mathfrak{a}}$ its image in $\bar{\hat{E}}_{\mathfrak{h}}$, we have $\pi(a_{u,\delta}\bar{\mathfrak{a}}a_{u,\delta}^{-1})={{}^{u}}\pi(\bar{\mathfrak{a}})$, which is just $\pi(\bar{\mathfrak{a}})$ by the assumption of $u$. Therefore, we have $a_{u,\delta}\bar{\mathfrak{a}}a_{u,\delta}^{-1}=\lambda_{\mathfrak{a}}\bar{\mathfrak{a}}$ for some $\lambda_{\mathfrak{a}}\in k^{*}$. Furthermore, it is clear that some $p$-th power of $a_{u,\delta}$ has to be in $(\mathbb{A}_{\delta}^{D})^{*}$, which is isomorphic to $k^{*}\times(j+J(\mathbb{A}_{\delta})^{D})$. Then this implies some $p$-th power of $\lambda_{\mathfrak{a}}$ being $1$ and hence $\lambda_{\mathfrak{a}}$ itself is trivial. Therefore, we have $a_{u,\delta}\hat{E}_{\mathfrak{h}}a_{u,\delta}^{-1}=\hat{E}_{\mathfrak{h}}$. In particular, $a_{u,\delta}E_{\mathfrak{h}}a_{u,\delta}^{-1}$ is also a complement of $j+J(\mathbb{A}_{\delta})^{D}$ in $\hat{E}_{\mathfrak{h}}$.
Since $E_{\mathfrak{h}}$ is a $p^{\prime}$-group, by [29, Lemma (45.6)], $a_{u,\delta}E_{\mathfrak{h}}a_{u,\delta}^{-1}$ and $E_{\mathfrak{h}}$ are conjugate to each other by an element of $j+J(\mathbb{A}_{\delta})^{D}$. Up to some adjustment, we can assume that $a_{u,\delta}E_{\mathfrak{h}}a_{u,\delta}^{-1}=E_{\mathfrak{h}}$. For any $\mathfrak{a}\in E_{\mathfrak{h}}$, it is clear that $\pi(a_{u,\delta}\mathfrak{a}a_{u,\delta}^{-1})=\pi(\mathfrak{a})$. Then we have $a_{u,\delta}\mathfrak{a}a_{u,\delta}^{-1}=\mathfrak{a}$. Then the lemma holds. ∎
5.19.
Fix an element $u$ of $P$ satisfying that it acts trivially on $E_{\mathfrak{h}}$ through the canonical map $P\longrightarrow P/C_{P}(D)$. By Lemma 5.18, we can assume that $a_{u,\delta}$ acts trivially on the subgroup $E_{\mathfrak{h}}$ of $N_{\mathbb{A}_{\delta}}(D)$ through the conjugation action. Then we have the twisted $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule ${{}_{(u,u)}}M_{\mathrm{st}}$, which is equal to $M_{\mathrm{st}}$ as $\mathcal{O}$-modules, with $(uau^{-1},uxu^{-1})$ acting on $m\in{{}_{(u,u)}}M_{\mathrm{st}}$ as $(a,x)$ on $m\in M_{\mathrm{st}}$.
@@
Keep the notation and assumption as above. Then the two modules $M_{\mathrm{st}}$ and ${{}_{(u,u)}}M_{\mathrm{st}}$ are isomorphic to each other as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules.
@@
Recall that from the paragraph 5.9 that we have an element $s_{u}$ of $S=\mathrm{End}_{\mathcal{O}}(V)$ such that $s_{u}(x\cdot\mathrm{v})=(uxu^{-1})\cdot s_{u}(\mathrm{v})$ for any $x\in L$ and any $\mathrm{v}\in V$. At the same time, we have a unitary subalgebra $\mathop{\bigoplus}\limits_{\tilde{g}\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{e}_{\tilde{g}}$ of $\mathbb{A}_{\delta}$ isomorphic to $\mbox{$\mathcal{O}$}L$ (see the proof of Proposition 4.12). By Lemma 5.18, we can assume that $a_{u,\delta}\mathfrak{e}_{\tilde{g}}a_{u,\delta}^{-1}=\mathfrak{e}_{\tilde{g}}$ for any $\tilde{g}\in E_{\mathfrak{h}}$. Note that $M_{\mathrm{st}}=\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((V\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cdot e)$. We can define the following map from $M_{\mathrm{st}}$ to ${{}_{(u,u)}}M_{\mathrm{st}}$
$\tau:M_{\mathrm{st}}\longrightarrow{{}_{(u,u)}}M_{\mathrm{st}},~~~~\mathfrak{a}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((\mathrm{v}\mathop{\otimes}\limits_{\mathcal{O}}\mathfrak{a}_{\delta})\cdot e)\mapsto u\mathfrak{a}a_{u,\delta}^{-1}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((s_{u}(\mathrm{v})\mathop{\otimes}\limits_{\mathcal{O}}a_{u,\delta}\mathfrak{a}_{\delta}a_{u,\delta}^{-1})\cdot e),$
where $\mathfrak{a}$ is in $\mathbb{A}$ and $\mathrm{v}$ is in $V$ and $\mathfrak{a}_{\delta}$ is in $\mathbb{A}_{\delta}$. It is easy to check that $\tau$ is well-defined and a bijection. Let $\mathfrak{a}^{\prime}$ and $w$ and $\mathfrak{e}_{\tilde{g}}$ be in $\mathbb{A}$ and $D$ and $E_{\mathfrak{h}}$, respectively.
@@
$=$~ $u\mathfrak{a}^{\prime}\mathfrak{a}a_{u,\delta}^{-1}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((s_{u}((w\tilde{g})^{-1}(\mathrm{v}))\mathop{\otimes}\limits_{\mathcal{O}}a_{u,\delta}\mathfrak{a}_{\delta}\cdot w\mathfrak{e}_{\tilde{g}}a_{u,\delta}^{-1})\cdot e)$
$=$~ $(u\mathfrak{a}^{\prime}u^{-1})\cdot u\mathfrak{a}a_{u,\delta}^{-1}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((uw\tilde{g}u^{-1})^{-1}\cdot s_{u}(\mathrm{v})\mathop{\otimes}\limits_{\mathcal{O}}a_{u,\delta}\mathfrak{a}_{\delta}a_{u,\delta}^{-1}\cdot(a_{u,\delta}w\mathfrak{e}_{\tilde{g}}a_{u,\delta}^{-1})\cdot e)$
@@
where the multiplications in the last equation is taken in ${{}_{(u,u)}}M_{\mathrm{st}}$. We complete the proof of this proposition. ∎
Corollary 5.21.
(i) Let $u$ be an element of $C_{P}(D)$. Then the two modules $M_{\mathrm{st}}$ and ${{}_{(u,u)}}M_{\mathrm{st}}$ are isomorphic to each other as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules.
(ii) Suppose that $N_{G}(P_{\gamma})$ controls fusion of the block $b$. Then the two modules $M_{\mathrm{st}}$ and ${{}_{(u,u)}}M_{\mathrm{st}}$ are isomorphic to each other as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules for any $u\in P$.
Proof.
Statement (i) follows from Proposition 5.18 directly. Now we assume that $N_{G}(P_{\gamma})$ controls fuison of the block $b$. Let $(P,\bar{b}_{P})$ and $(D,\bar{b}_{D})$ be the $b$-Brauer pairs associated with $P_{\gamma}$ and $D_{\delta}$, respectively. By [25, Proposition 3.3], the block $\bar{b}_{D}$ is nilpotent and so $N_{G}(D,\bar{b}_{D})=N_{G}(D_{\delta})$. Since $N_{G}(P_{\gamma})$ controls fuison of the block $b$, we have
$N_{G}(D_{\delta})=N_{G}(D,\bar{b}_{D})=N_{G}(P,\bar{b}_{D})C_{G}(D)=N_{G}(P_{\gamma})C_{G}(D)$
due to the well-known fact that $N_{G}(P,\bar{b}_{D})=N_{G}(P_{\gamma})$. As a block of $PC_{G}(D)$, it is clear that $b_{D}$ is still nilpotent with a defect group $P$. Therefore, $(P,\bar{b}_{P})$ is still a maximal $b_{D}$-Brauer pair of the block $b_{D}$ of $PC_{G}(D)$. Then
$N_{G}(P_{\gamma})\cap C_{G}(D)=$ $~N_{G}(P,\bar{b}_{P})\cap C_{G}(D)\cap PC_{G}(D)$
$=$ $~N_{PC_{G}(D)}(P,\bar{b}_{D})\cap C_{G}(D)$
$=$ $~PC_{G}(P)\cap C_{G}(D)$
$=$ $~C_{P}(D)C_{G}(P).$
Here, the third equation is due to the definition of nilpotent blocks. Hence, $N_{G}(D_{\delta})/C_{G}(D)$ is isomorphic to $N_{G}(P_{\gamma})/C_{P}(D)C_{G}(P)$ and this isomorphism is just induced by restriction. Then we can identify $E_{\mathfrak{h}}$ with the inertial quotient $E=N_{G}(P_{\gamma})/PC_{G}(P)$ of the block $b$.
On the other hand, by [11, Lemma 4.28], we have $P=[P,E]\cdot C_{P}(E)$, where $[P,E]$ is the commutator subgroup of $P$ and $E$. By the definition of the hyperfocal subgroup, it is clear that $[P,E]$ is contained in $D$. By the assumption that the block $b$ is a hyperfocal abelian Frobenius block, $D\cap C_{P}(E)=C_{D}(E)$ is trivial. Therefore, we have $D=[P,E]$ and $P=D\rtimes C_{P}(E)$. Thus every element $u$ of $P$ satisfies that it acts trivially on $E_{\mathfrak{h}}$ through the canonical map $P\longrightarrow P/C_{P}(D)$. Now statement (ii) follows from Proposition 5.18. ∎
@@ 6.1 Klein four group case
Keep the notation from the last three sections. In this subsection, we will assume that the prime number $p$ is equal to $2$ and the hyperfocal subgroup $D$ is a Klein four group. In this case the hyperfocal quotient inertial $E_{\mathfrak{h}}$ is a cyclic group of order $3$ and acts freely on $D-\{1\}$. Then we can apply Theorem 5.14 to this case and get a stable equivalence of Morita type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$. Here $L=D\rtimes E_{\mathfrak{h}}$ which is just $\mathrm{A}_{4}$ in this section. Here, $\mathrm{A}_{n}$ denotes the alternating group on $n$ letters for any positive integer $n$. In fact, we can classify the Morita equivalent classes and the derived equivalent classes of the hyperfocal subalgebras in this case, which is similar to the case where the defect group is a Klein four group.
@@
Denote $\mbox{$\mathcal{K}$}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}$ and $k\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}$ by $\hat{\mathbb{A}}$ and $\bar{\mathbb{A}}$, respectively. We have analogue notation for the source algebra $A$. It is clear that $\hat{\mathbb{A}}$ is a semi-simple $\mathcal{K}$-algebra (see [13]). By [3, (29.20)], there is a finite extension $\mbox{$\mathcal{K}$}^{\prime}$ of $\mathcal{K}$ such that the semi-simple $\mbox{$\mathcal{K}$}^{\prime}$-algebra $\mbox{$\mathcal{K}$}^{\prime}\mathop{\otimes}\limits_{\mathcal{K}}\hat{\mathbb{A}}$ is split. Since $k$ is algebraically closed, the field $\mbox{$\mathcal{K}$}^{\prime}$ determines a complete $p$-modular system $(\mbox{$\mathcal{K}$}^{\prime},\mbox{$\mathcal{O}$}^{\prime},k)$ such that $\mbox{$\mathcal{O}$}^{\prime}$ is an extension of $\mathcal{O}$ by [28, Chapter 2,§2, Proposition 3]. Replacing $\mathcal{O}$ by $\mbox{$\mathcal{O}$}^{\prime}$, we can assume that the $\mathcal{K}$-algebra $\hat{\mathbb{A}}$ is a split semi-simple $\mathcal{K}$-algebra. Then we can borrow the notation from the paragraphs 2.7-2.8 for these three $\mathcal{O}$-algebras $A$ and $\mathbb{A}$ and $\mathbb{A}_{\delta}$.
@@
By Theorem 5.14, there is a stable equivalence of Morita type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$. It is clear that the Cartan matrix of $\mbox{$\mathcal{O}$}L$ is nonsingular. Then by [16, Proposition 4.14.13], the Cartan matrix of $\bar{\mathbb{A}}$ is also nonsingular. Note that the hyperfocal subalgebra $\mathbb{A}$ is a symmetric $\mathcal{O}$-algebra (see [10, Theorem 1.1]). Then by [12, Proposition 3.1] the stable equivalence between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$ induces an isometry between $L^{0}(\mathbb{A})$ and $L^{0}(\mbox{$\mathcal{O}$}L)$. It is clear that $L^{0}(\mbox{$\mathcal{O}$}L)$ is a free $\mathbb{Z}$-module of rank one and setting $\zeta$ to be a generator, $\langle\zeta,\zeta\rangle_{\mathcal{O}L}=4$. Then $L^{0}(\mathbb{A})$ is also a free $\mathbb{Z}$-module of rank one which implies $\mathrm{k}(\mathbb{A})-l(\mathbb{A})=1$. Suppose that $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})=\{U_{1},U_{2},\cdots,U_{t}\}$ and $\sum_{a=1}^{t}n_{a}[U_{a}]\in L^{0}(\mathbb{A})$ is a generator, where $n_{1},n_{2},\cdots,n_{t}$ are integers. Then we have $\sum\limits_{a=1}^{t}n_{a}^{2}=4$. The only possibility is that $t=4$ and $n_{a}=\pm 1$ for $a=1,2,3,4$. We are done. ∎
6.3.
@@
Let $\{T_{1},T_{2},T_{3}\}$ be a set of representatives of the isomorphism classes of simple $kL$-modules. Denote by $\Omega=\Omega_{kL}$ the Heller operator of $kL$. By [17, Corollary 7.2.11], there are six isomorphism classes of indecomposable $kL$-modules of dimension $2$, denoted by $T_{\mathrm{i}}^{\mathrm{j}}$ with distinct $\mathrm{i}$ and $\mathrm{j}$ in $\{1,2,3\}$. Furthermore, the indecomposable $kL$-module $T_{\mathrm{i}}^{\mathrm{j}}$ has composition series $T_{\mathrm{j}}$, $T_{\mathrm{i}}$ and $\Omega(T_{\mathrm{i}}^{\mathrm{j}})=T_{\mathrm{j}}^{\mathrm{l}}$, where $\mathrm{l}\in\{1,2,3\}-\{\mathrm{i,\mathrm{j}}\}$. Since the two algebras $\mathbb{A}$ and $\mathbb{A}_{\delta}$ are Morita equivalent, by Lemma 6.2, we can set $\{\bar{W}_{1},\bar{W}_{2},\bar{W}_{3}\}$ to be a set of representatives of the isomorphism classes of simple $\bar{\mathbb{A}}_{\delta}$-modules. Now we can state the following main theorem in this section and give a proof of it, which is similar to the one of [17, Theorem 12.1.2].

### v2

@@ 5 A stable equivalence of Morita type
Recall that we denote $N_{G}(D_{\delta})/C_{G}(D)$ by $\tilde{N}$. By the paragraph 2.12, we can get a short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${k^{\times}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{A_{\delta}}(D)/(j+J(A_{\delta}^{D}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\tilde{N}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
since $D$ is abelian. Denote by $\pi$ the sujective homomorphism of the above sequence. Note that $\tilde{N}$ is a $p$-nilpotent group with Sylow $p$-subgroup $\tilde{P}$ and normal $p$-complement $E_{\mathfrak{h}}$. Here, $\tilde{P}$ denotes $PC_{G}(D)/C_{G}(D)$ which is isomorphic to $P/C_{P}(D)$. Obviously, the second cohomology group $H^{2}(\tilde{P};k^{\times})$ of $\tilde{P}$ is trivial. Since $D\rtimes E_{\mathfrak{h}}$ is a Frobenius group, the second cohomology group $H^{2}(E_{\mathfrak{h}};k^{\times})$ of $E_{\mathfrak{h}}$ is also trivial. In conclusion, the second cohomology group $H^{2}(\tilde{N};k^{\times})$ of $\tilde{N}$ is trivial. This means that the short exact sequence above splits. So there is a subgroup $\hat{F}$ of $N_{A_{\delta}}(D)$ such that it contains $j+J(A_{\delta}^{D})$ and intersects with $k^{\times}$ trivially and $\hat{F}/(j+J(A_{\delta}^{D}))$ is isomorphic to $\tilde{N}$. So for any $v\in P$, the element $a_{v,\delta}$ can be chosen such that the set $P(j+J(A_{\delta}^{D}))=\{a_{v,\delta}x\mid v\in P,x\in j+J(A_{\delta}^{D})\}$ is a Sylow $p$-subgroup of $\hat{F}/(j+J(A_{\delta}^{D}))$. Denote by $\hat{F}_{\mathfrak{h}}$ a normal subgroup of $\hat{F}$ containing $j+J(A_{\delta}^{D})$ such that $\hat{F}_{\mathfrak{h}}/(j+J(A_{\delta}^{D}))$ is a normal $p$-complement of $\hat{F}/(j+J(A_{\delta}^{D}))$. In particular, ${{}^{a_{v,\delta}}}\hat{F}_{\mathfrak{h}}=\hat{F}_{\mathfrak{h}}$ for any $v\in P$ and $\hat{F}_{\mathfrak{h}}/(j+J(A_{\delta}^{D}))$ is isomorphic to $E_{\mathfrak{h}}$. Hence, this yields another short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${j+J(A_{\delta}^{D})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\hat{F}_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1,}$
which also splits by [29, Lemma (45.6)] since $E_{\mathfrak{h}}$ is a $p^{\prime}$-group. So there is a subgroup $\mathfrak{F}_{\mathfrak{h}}=\{{\rm f}_{\tilde{g}}\mid\tilde{g}\in E_{\mathfrak{h}}\}$ of $\hat{F}_{\mathfrak{h}}$ such that it is isomorphic to $E_{\mathfrak{h}}$ and intersects with $j+J(A_{\delta}^{D})$ trivially.
[…]
For the remainder of this paper, we can assume that the element $a_{u,\delta}$ in $N_{A_{\delta}}(D)$ satisfies the properties in Lemma 5.18 and $a_{u,\delta}(wj)a_{u,\delta}^{-1}=(uwu^{-1})j$ for any $u\in P$ and $w\in D$. Now we can state the $\Delta(P)$-equivariant property of the bimodule $M_{{\rm st}}$.
@@
Keep the notation. Then the two modules $M_{\mathrm{st}}$ and ${{}_{(u,u)}}M_{\mathrm{st}}$ are isomorphic to each other as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules.
@@
Recall that from the paragraph 5.9 that we have an invertible element $s_{u}$ of $S=\mathrm{End}_{\mathcal{O}}(V)$ such that $s_{u}(x\cdot\mathrm{v})=(uxu^{-1})\cdot s_{u}(\mathrm{v})$ for any $x\in L$ and any $\mathrm{v}\in V$. Note that $M_{\mathrm{st}}=\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((V\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cdot e)$. We can define the following map from $M_{\mathrm{st}}$ to ${{}_{(u,u)}}M_{\mathrm{st}}$
$\tau:M_{\mathrm{st}}\longrightarrow{{}_{(u,u)}}M_{\mathrm{st}},~~~~\mathfrak{a}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((\mathrm{v}\mathop{\otimes}\limits_{\mathcal{O}}\mathfrak{a}_{\delta})\cdot e)\mapsto u^{-1}\mathfrak{a}a_{u,\delta}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((s_{u}^{-1}(\mathrm{v})\mathop{\otimes}\limits_{\mathcal{O}}a_{u,\delta}^{-1}\mathfrak{a}_{\delta}a_{u,\delta})\cdot e),$
where $\mathfrak{a}$ is in $\mathbb{A}$ and $\mathrm{v}$ is in $V$ and $\mathfrak{a}_{\delta}$ is in $\mathbb{A}_{\delta}$. It is easy to check that $\tau$ is well-defined and a bijection. Let $\mathfrak{a}^{\prime}$ and $w$ and $\tilde{g}$ be in $\mathbb{A}$ and $D$ and $E_{\mathfrak{h}}$, respectively.
@@
$=$~ $u^{-1}\mathfrak{a}^{\prime}\mathfrak{a}a_{u,\delta}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((s_{u}^{-1}((w\tilde{g})^{-1}(\mathrm{v}))\mathop{\otimes}\limits_{\mathcal{O}}a_{u,\delta}^{-1}\mathfrak{a}_{\delta}\cdot w\mathfrak{e}_{\tilde{g}}a_{u,\delta})\cdot e)$
$=$~ $(u^{-1}\mathfrak{a}^{\prime}u)\cdot u^{-1}\mathfrak{a}a_{u,\delta}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((u^{-1}w\tilde{g}u)^{-1}\cdot s_{u}^{-1}(\mathrm{v})\mathop{\otimes}\limits_{\mathcal{O}}a_{u,\delta}^{-1}\mathfrak{a}_{\delta}a_{u,\delta}\cdot(a_{u,\delta}^{-1}w\mathfrak{e}_{\tilde{g}}a_{u,\delta})\cdot e)$
@@
where the multiplications in the last equation are taken in ${{}_{(u,u)}}M_{\mathrm{st}}$. We complete the proof of this proposition. ∎
@@ 6.1 Klein four group case
Keep the notation from the last three sections. In this subsection, we will assume that the prime $p$ is equal to $2$ and the hyperfocal subgroup $D$ is a Klein four group. In this case the hyperfocal quotient inertial $E_{\mathfrak{h}}$ is a cyclic group of order $3$ and acts freely on $D-\{1\}$. Then we can apply Theorem 5.14 to this case and get a stable equivalence of Morita type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$. Here $L=D\rtimes E_{\mathfrak{h}}$ which is just $\mathrm{A}_{4}$ in this section. Here, $\mathrm{A}_{n}$ denotes the alternating group on $n$ letters for any positive integer $n$. In fact, we can classify the Morita equivalence classes and the derived equivalence classes of the hyperfocal subalgebras in this case, which is similar to the case where the defect group is a Klein four group.
@@
Denote $\mbox{$\mathcal{K}$}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}$ and $k\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}$ by $\hat{\mathbb{A}}$ and $\bar{\mathbb{A}}$, respectively. We use the analogous notation for the source algebra $A$. It is clear that $\hat{\mathbb{A}}$ is a semi-simple $\mathcal{K}$-algebra (see [12]). By [3, (29.20)], there is a finite extension $\mbox{$\mathcal{K}$}^{\prime}$ of $\mathcal{K}$ such that the semi-simple $\mbox{$\mathcal{K}$}^{\prime}$-algebra $\mbox{$\mathcal{K}$}^{\prime}\mathop{\otimes}\limits_{\mathcal{K}}\hat{\mathbb{A}}$ is split. Since $k$ is algebraically closed, the field $\mbox{$\mathcal{K}$}^{\prime}$ determines a complete $p$-modular system $(\mbox{$\mathcal{K}$}^{\prime},\mbox{$\mathcal{O}$}^{\prime},k)$ such that $\mbox{$\mathcal{O}$}^{\prime}$ is an extension of $\mathcal{O}$ by [28, Chapter 2, §2, Proposition 3]. Replacing $\mathcal{O}$ by $\mbox{$\mathcal{O}$}^{\prime}$, we can assume that the $\mathcal{K}$-algebra $\hat{\mathbb{A}}$ is a split semi-simple $\mathcal{K}$-algebra. Then we can borrow the notation from the paragraphs 2.6-2.7 for these three $\mathcal{O}$-algebras $A$ and $\mathbb{A}$ and $\mathbb{A}_{\delta}$.
@@
By Theorem 5.14, there is a stable equivalence of Morita type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$. By Corollary 5.15, the Cartan matrix of $\bar{\mathbb{A}}$ is nonsingular. Note that the hyperfocal subalgebra $\mathbb{A}$ is a symmetric $\mathcal{O}$-algebra (see [10, Theorem 1.1]). Then by [11, Proposition 3.1] the stable equivalence between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$ induces an isometry between $L^{0}(\mathbb{A})$ and $L^{0}(\mbox{$\mathcal{O}$}L)$. It is clear that $L^{0}(\mbox{$\mathcal{O}$}L)$ is a free $\mathbb{Z}$-module of rank one. Set $\zeta$ to be a generator. It is well-known that $\langle\zeta,\zeta\rangle_{\mathcal{O}L}=4$. Then $L^{0}(\mathbb{A})$ is also a free $\mathbb{Z}$-module of rank one which implies that $\mathrm{k}(\mathbb{A})-l(\mathbb{A})=1$. Suppose that $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})=\{\hat{U}_{1},\hat{U}_{2},\cdots,\hat{U}_{t+1}\}$ and $\mathrm{Irr}_{k}(\mathbb{A})=\{\bar{M}_{1},\bar{M}_{2},\cdots,\bar{M}_{t}\}$ for some positive integer $t$.
By [7, Theorem 1.1], $l(A)$ is equal to $2$ or $3$. Suppose that $l(A)=3$. In this case, we have $P=C_{P}(D)$ by [7, Theorem 1.1 (i) and Proposition 2.3 (a)]. Then by Corollary 4.10, every simple $\bar{\mathbb{A}}$-module is $P$-stable. In particular, every simple $\bar{\mathbb{A}}$-module can be uniquely extended to a simple $\bar{A}$-module. This implies that $\bar{A}$ and $\bar{\mathbb{A}}$ have the same number of isomorphism classes of simple modules. In particular, we have $l(\mathbb{A})=3$ and then ${\rm k}(\mathbb{A})=4$.
Now suppose that $l(A)$ is equal to $2$. In this case, the quotient group $P/C_{P}(D)$ is a group of order $2$ by [7, Theorem 1.1 (ii) and Proposition 2.3 (b)]. Then by Corollary 4.10 again, the quotient group $P/C_{P}(D)$ can act on the set ${\rm Irr}_{k}(\mathbb{A})$. By Clifford Theory, $l(\mathbb{A})$ has three possibilities: $2$, or $3$, or $4$. Assume that $l(\mathbb{A})$ equals $2$. Then ${\rm k}(\mathbb{A})$ equals $3$. Set $\sum\limits_{a=1}^{3}n_{a}[\hat{U}_{a}]$ is a generator of $L^{0}(\mathbb{A})$, where $n_{1}$ and $n_{2}$ and $n_{3}$ are integers. Since there is an isometry between $L^{0}(\mathbb{A})$ and $L^{0}(\mbox{$\mathcal{O}$}L)$, we have
$\langle\sum\limits_{a=1}^{3}n_{a}[\hat{U}_{a}],\sum\limits_{a=1}^{3}n_{a}[\hat{U}_{a}]\rangle_{\mathbb{A}}=\sum\limits_{a=1}^{3}n_{a}^{2}=4.$
It is easy to see that all but one of the coefficients are zero and then $2[\hat{U}_{a}]$ belongs to $L^{0}(\mathbb{A})$ for some $1\leq a\leq 3$. This is impossible. Assume that $l(\mathbb{A})$ equals $4$. In this subcase, without loss of generality, we can assume that $\bar{M}_{1}$ and $\bar{M}_{2}$ are permuted by $P/C_{P}(D)$ and $\bar{M}_{3}$ and $\bar{M}_{4}$ are permuted by $P/C_{P}(D)$, respectively. Set $\mathbb{A}_{C_{P}(D)}$ to be $\mathop{\bigoplus}\limits_{uD\in C_{P}(D)/D}\mathbb{A}u$. Then for any $1\leq a\leq 4$, the simple $\bar{\mathbb{A}}$-module $\bar{M}_{a}$ can be uniquely extended to a simple $\bar{\mathbb{A}}_{C_{P}(D)}$-module, which we still denote by $\bar{M}_{a}$. Here, $\bar{\mathbb{A}}_{C_{P}(D)}$ is the $k$-algebra $k\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{C_{P}(D)}$. Therefore, we can get that ${\rm Irr}_{k}(A)=\{\bar{A}\mathop{\otimes}\limits_{\bar{\mathbb{A}}_{C_{P}(D)}}\bar{M}_{1},\bar{A}\mathop{\otimes}\limits_{\bar{\mathbb{A}}_{C_{P}(D)}}\bar{M}_{3}\}$. Note that ${\rm dim}_{k}(\bar{A}\mathop{\otimes}\limits_{\bar{\mathbb{A}}_{C_{P}(D)}}\bar{M}_{1})=2{\rm dim}_{k}(\bar{M}_{1})$ and ${\rm dim}_{k}(\bar{A}\mathop{\otimes}\limits_{\bar{\mathbb{A}}_{C_{P}(D)}}\bar{M}_{3})=2{\rm dim}_{k}(\bar{M}_{3})$. This means that all simple $\bar{A}$-modules have even dimensions. But by [29, Proposition (44.9)], there exists some simple $\bar{A}$-module of odd dimension, a contradiction.
In conclusion, we have $l(\mathbb{A})=3$ and then ${\rm k}(\mathbb{A})=4$.
Remark 6.3.
Another possible way to calculate ${\rm k}(\mathbb{A})$ is to analyse the coefficients of a generator $\sum\limits_{a=1}^{t+1}n_{a}[\hat{U}_{a}]$ of $L^{0}(\mathbb{A})$ through the equations $\langle\sum\limits_{a=1}^{t+1}n_{a}[\hat{U}_{a}],\sum\limits_{a=1}^{t+1}n_{a}[\hat{U}_{a}]\rangle_{\mathbb{A}}=\sum\limits_{a=1}^{t+1}n_{a}^{2}=4$. This is exactly the approach taken when calculating $|{\rm Irr}(b)|$ for the block $b$ with a Klein four defect group (see [16, Corollary 12.2.5]). This way is easy and efficient for the block due to a well-known fact of the block that an irreducible ordinary character is a $\mathbb{Z}$-linear combination of some characters of projective modules if and only if it belongs to a block of a trivial defect group. It is reasonable to speculate that a simple $\hat{\mathbb{A}}-$module $\hat{U}$ belongs to ${\rm Pr}_{\mathcal{O}}(\mathbb{A})$ if and only if the hyperfocal subgroup is trivial for the hyperfocal subalgebra $\mathbb{A}$ in general. But this seems not easy to prove for us.
6.4.
@@
Let $\{T_{1},T_{2},T_{3}\}$ be a set of representatives of the isomorphism classes of simple $kL$-modules. Moreover, we can set $T_{1}$ to be the trivial simple $kL$-module $k$. Denote by $\Omega=\Omega_{kL}$ the Heller operator of $kL$. By [16, Corollary 7.2.11], there are six isomorphism classes of indecomposable $kL$-modules of dimension $2$, denoted by $T_{\mathrm{i}}^{\mathrm{j}}$ with distinct $\mathrm{i}$ and $\mathrm{j}$ in $\{1,2,3\}$. Furthermore, the indecomposable $kL$-module $T_{\mathrm{i}}^{\mathrm{j}}$ has composition series $T_{\mathrm{j}}$, $T_{\mathrm{i}}$ and $\Omega(T_{\mathrm{i}}^{\mathrm{j}})=T_{\mathrm{j}}^{\mathrm{l}}$, where $\mathrm{l}\in\{1,2,3\}-\{\mathrm{i,\mathrm{j}}\}$. Since the two algebras $\mathbb{A}$ and $\mathbb{A}_{\delta}$ are Morita equivalent, by Lemma 6.2, we can set $\{\bar{W}_{1},\bar{W}_{2},\bar{W}_{3}\}$ to be a set of representatives of the isomorphism classes of simple $\bar{\mathbb{A}}_{\delta}$-modules. Now we can state the following main theorem in this subsection, which yields the Morita equivalence classes and the derived equivalence classes of the hyperfocal subalgebra $\mathbb{A}$ with a Klein four group.